40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Dec 4 23:16:35 2007, tobin, Update, PSL, ISS 
    Reply  Wed Dec 5 17:49:57 2007, tobin, Update, PSL, ISS rin.pdf
       Reply  Wed Dec 5 18:22:03 2007, tobin, Update, PSL, ISS dark noise sense_noise.pdf
          Reply  Wed Dec 5 20:32:51 2007, tobin, Update, PSL, ISS dark noise 
             Reply  Thu Dec 6 19:30:43 2007, tobin, Update, PSL, ISS dark noise - 60 Hz! iss.pdf
                Reply  Fri Dec 7 19:14:30 2007, tobin, Update, PSL, ISS dark noise - ground loop enlightenment 
Message ID: 169     Entry time: Wed Dec 5 18:22:03 2007     In reply to: 167     Reply to this: 171
Author: tobin 
Type: Update 
Category: PSL 
Subject: ISS dark noise 
Attached is a plot of the dark noise spectrum of the ISS photodiodes (1) before fooling with them, (2) after replacing the 4151's with 4131's (improvement!), and (3) after replacing the cables and changing the wiring (disaster!).
Attachment 1: sense_noise.pdf  29 kB  | Hide | Hide all
sense_noise.pdf
ELOG V3.1.3-